OpenStudy (anonymous):

Integration problem to calculate surface area of a sphere by rotating the half-circle. Equation to integrate and entire problem statement in next box.

OpenStudy (anonymous):

\[S = 2 \pi \int\limits \sqrt{R^2-x^2} \ \sqrt{1 + \frac{x^2}{1 - x^2}} \ dx\]

OpenStudy (anonymous):

R is a constant.

OpenStudy (amistre64):

\[\int \sqrt{\frac{R^2-x^2}{1-x^2}}dx\]is what i get to by simplifying ... not that it helps :)

OpenStudy (anonymous):

Right, I forget about that step!!

OpenStudy (anonymous):

forgot

OpenStudy (amistre64):

\[y' = \frac{R-x}{y}\] \[(y')^2 = \frac{(R-x)^2}{y2}=\left( \frac{R-x}{R^2-x^2}\right)^2\]trying to see if this is a "given" as stated

OpenStudy (amistre64):

ack, i treated R as a variable lol

OpenStudy (amistre64):

still y^2 = R^2-x^2; not 1-x^2

OpenStudy (anonymous):

Thanks for any help. I'll think about it today. Funny part is I know the answer has to be Rx.

OpenStudy (amistre64):

you used x^2+y^2 = 1 in the second general part there; when the Radius is spose to be R

OpenStudy (amistre64):

\[\int \sqrt{\frac{R^2-x^2}{(\cancel{1})R^2-x^2}}dx\]

OpenStudy (anonymous):

Not following. The surface element does not contain R, only the first term which is f(x) and that has R^2.

OpenStudy (anonymous):

OHHH

OpenStudy (anonymous):

I subbed for y in ds

OpenStudy (amistre64):

the circle is: x^2 + y^2 = R^2 y = sqrt(R^2 - x^2) since y' = -x/y, and (y')^2 = .....

OpenStudy (anonymous):

Thank you!

OpenStudy (amistre64):

youre welcome

OpenStudy (amistre64):

\[\int \sqrt{\frac{R^2(R^2-x^2)}{R^2-x^2}}dx=2piR\int_{-R}^{R} dx\]